# Liquid Resistance Liquid Capacitance Networks

### v2

@@ Abstract
We introduce liquid-resistance liquid-capacitance neural networks (LRCs), a neural-ODE model which considerably improves the smoothness, accuracy, and biological plausibility of electrical equivalent circuits (EECs), liquid time-constant networks (LTCs), and saturated liquid time-constant networks (STCs), respectively. We also introduce LRC units (LRCUs), as a very efficient and accurate gated RNN-model, which results from solving LRCs with an explicit Euler scheme using just one unfolding. We empirically show and formally prove that the liquid capacitance of LRCs considerably dampens the oscillations of LTCs and STCs, while at the same time dramatically increasing accuracy even for cheap solvers. We experimentally demonstrate that LRCs are a highly competitive alternative to popular neural ODEs and gated RNNs in terms of accuracy, efficiency, and interpretability, on classic time-series benchmarks and a complex autonomous-driving lane-keeping task.
@@ 1 Introduction
Electrical equivalent circuits (EECs) are the foremost mathematical model used in neuroscience for capturing the dynamic behavior of biological neurons (Kandel et al., 2000; Wicks et al., 1996). EECs associate the membrane of a neuron with a capacitor, whose potential varies according to the sum of the stimulus, leaking, and synaptic currents, passing through the membrane, respectively. As a consequence, EECs are formulated as a set of capacitor ordinary differential equations (ODEs).
@@
EECs however, received little attention in the ML community, due to their ODE nature, except for continuous-time recurrent neural networks (CT-RNNs) (Funahashi and Nakamura, 1993), which are arguably EECs with electrical synapses (Farsang et al., 2023). This changed with the advent of Neural ODEs (Chen et al., 2018; Rubanova et al., 2019; Dupont et al., 2019), and EECs with chemical synapses were called liquid time-constant networks (LTCs) in Lechner et al. (2019); Hasani et al. (2020); Lechner et al. (2020), as their time constant depends on both the state and the input.
@@
The main promise of LTCs and Neural ODEs is to better capture physical reality and bridge the gap between natural sciences and ML. They were shown to be more interpretable (Lechner et al., 2020; Farsang et al., 2023) and able to recover missing data in irregularly-sampled time series (Chen et al., 2018; Rubanova et al., 2019; Dupont et al., 2019; Brouwer et al., 2019; Kidger et al., 2020; li et al., 2023). An important obstacle in the use of Neural ODEs and especially LTCs however, is their stiff oscillatory nature, which requires the use of expensive ODE solvers (Biloš et al., 2021). A best solver moreover, cannot fully remove oscillations and improve accuracy, as shown in Figures 1 and 2.
@@
In this paper, we show that the stiff dynamic behavior of LTCs results from abstraction. For simplicity, EECs and LTCs first ignore the saturation aspects of membrane’s ionic channels (Alberts et al., 2008), as discussed in Farsang et al. (2023). To take such aspects into account, they introduce saturated LTCs (called STCs). Second, and more importantly, EECs, LTCs and STCs also assume that membranal capacitance is constant, which is disproved by recent results demonstrating a nonlinear dependence on the neural state and input (Howell et al., 2015; Severin et al., 2022; Kumar et al., 2023).
Figure 1: Dynamic behavior of the output neuron in the Lane-Keeping Task (Lechner et al., 2022). Left: In LTCs, this neuron has a very stiff oscillatory behavior, with a relatively large validation loss. Moreover, LTCs fail to converge with an explicit Euler integration scheme of order one and just one unfolding. Middle: STCs in contrast, convergence with an explicit Euler integration scheme of order one and just one unfolding. However, the output neuron still exhibits stiff oscillatory behavior. Right: In LRCs, this neuron has a very smooth behavior, with a very small validation loss, even when using an explicit Euler scheme of order one and just one unfolding (no overhead). For a fair comparison of their properties, we used essentially the same number of parameters in all three models.
Time Per Best Lipschitz
Model Epoch Epoch Constant
LTC H6 10.06 min $52\pm 06$ $90{\pm}23{\times}10^{-4}$
STC E1 03.97 min $50\pm 22$ $101{\pm}25{\times}10^{-4}$
STC H6 10.13 min $69\pm 10$ $89{\pm}23{\times}10^{-4}$
LRC-A E1 04.05 min $43\pm 08$ $47{\pm}08{\times}10^{-4}$
LRC-S E1 04.16 min $22\pm 03$ $43{\pm}11{\times}10^{-4}$
Figure 2: Training loss in the first and validation loss in the second figure of LTCs, STCs, and LRCs for the Lane-Keeping Task. LTCs and STCs use a constant, unit elastance, whereas LRCs use the liquid elastance of Equations (4-5). The results for LRCs solved with an explicit Euler integration scheme with one unfolding (called LRCUs) are substantially better, both in terms of convergence speed and accuracy. Per epoch and curves, the training took around 10 minutes for 6 unfoldings per step and 4 minutes for 1 unfolding per step, which is 2.5 times faster. As the curves of LTCs and STCs with the hybrid solver converge in around 50-70 epochs, and the LRC-S and LRC-A in around 20 and 40, respectively, the total speedup until the best epoch of the latter with respect to the former is 3-7.8 times, while also being a much more accurate. Note that solving STCs with Explicit Euler and one unfolding is computationally inexpensive, but its validation loss is not very satisfactory.
@@
Adding a liquid membrane capacitance to STCs, considerably stabilizes their dynamic behavior, while also enhancing their accuracy, as shown in Figures 1,2 and proved in Theorems 1,2. We call the resulting model a liquid-resistance, liquid-capacitance neural network (LRC), as the LRC’s liquid time-constant is now factored this way. Due to the smooth LRC behavior, an explicit Euler integration scheme with one unfolding, is often enough to obtain a very small error. We call LRCs solved with this scheme, LRC units (LRCUs). They closely resemble and improve on the performance of gated recurrent units (Hochreiter and Schmidhuber, 1997; Cho et al., 2014; Zhou et al., 2016). For irregularly sampled time series, however, one can use more expensive solvers with LRCs, too.
@@
We experimentally evaluated LRCs with respect to their accuracy and speed on the classic gated-RNN and Neural-ODE benchmarks, respectively, and with respect to their interpretability on a complex autonomous-driving lane-keeping task. Our results show that LRCs outperform LSTMs, GRUs, and MGUs on RNN benchmarks, are able to successfully solve Neural-ODE tasks, and at the same time, are much more interpretable than LSTMs, GRUs, and MGUs on the lane-keeping task.
@@
-- We introduce Liquid Resistance Liquid Capacitance Networks (LRCs), which considerably improve the smoothness, accuracy, and biological plausibility of EEC, LTC, and STC models, respectively.
@@
-- We prove that the liquid capacitance of LRCs, considerably dampens the oscillations in LTCs and STCs, while at the same time dramatically increasing their accuracy even for cheap solvers.
@@
The rest of this paper is organized as follows. In Section 2 we review EECs, LTCs, and STCs. In Section 3 we introduce LRCs and LRCUs and discuss and prove their properties. In Section 4 we provide our very encouraging experimental results. Finally, in Section 5 we draw our conclusions.
@@ 2.1 Liquid Time-Constant Networks (LTCs)
Definition. EECs are a simplified electrical model, defining the dynamic behavior of the membrane-potential (MP) of a post-synaptic neuron, as a function of the MP of its pre-synaptic neurons, for electrical or chemical synapses (Kandel et al., 2000; Wicks et al., 1996). EECs with chemical synapses only, are also known in the ML community as LTCs (Lechner et al., 2019; Hasani et al., 2020; Lechner et al., 2020). An LTC with $m$ neurons and $n$ inputs is defined as follows:
$\begin{array}[]{c}\dot{h}_{i}=-f_{i}\,h_{i}+u_{i}\,e_{li}\\[5.69054pt] f_{i}=\sum_{j=1}^{m+n}g_{ji}\sigma(a_{ji}y_{j}+b_{ji})+g_{li}\\[5.69054pt] u_{i}=\sum_{j=1}^{m+n}k_{ji}\sigma(a_{ji}y_{j}+b_{ji})+g_{li}\end{array}$ (1)
It states that the rate of change of the MP $h_{i}$ of neuron $i$, is the sum of its forget current $-f_{i}\,h_{i}$ and its update current $u_{i}\,e_{li}$. The forget conductance $f_{i}$ is the liquid time-constant reciprocal of $h_{i}$. It depends on $y\,{=}\,[h,x]$, the MPs $h$ of presynaptic neurons and inputs $x$. Here, $g_{ji}$ is the maximum conductance of the synaptic channels, $a_{ji},b_{ji}$ are parameters controlling the sigmoidal probability of these channels to be open, and $g_{li}$, $e_{li}$ are the membrane’s leaking (resting) conductance and potential, respectively. In the update conductance $u_{i}$, parameters $k_{ji}\,{=}\,g_{ji}e_{ji}{/}e_{li}$ take the sign of $e_{ji}{/}e_{li}$, where $e_{ji}$ is the synaptic channel’s reversal potential, that is, the MP at which there is no net ionic flow.
Smoothness. The dynamic behavior of the LTC output neuron in the Lane-Keeping Task, is shown in Figure 1(left). It has a very stiff oscillatory behavior, and it fails to converge with an explicit Euler integration with one unfolding. The validation loss is also relatively large, as shown in Figure 2.
@@ 2.2 Saturated Liquid Time-Constant Networks (STCs)
Definition. For simplicity, LTCs ignore saturation in synaptic channels (Alberts et al., 2008). One can take this aspect into account, by constraining the forget and signed-update conductances $f_{i}$ and $u_{i}$ with a sigmoid and a hyperbolic tangent, to range within [0,1] and [-1,1], respectively. Letting $f_{i}$ and $u_{i}$ be as above, this leads to the definition of saturated LTCs (STCs) of Farsang et al. (2023):
$\begin{array}[]{c}\dot{h}_{i}=-\sigma(f_{i})\,h_{i}+\tau(u_{i})\,e_{li}\end{array}$ (2)
Smoothness. The dynamic behavior of the STC output neuron in the Lane-Keeping Task is shown in Figure 1 (middle). It is smoother than in LTCs, and it succeeds to converge with an explicit Euler integration with one unfolding. However, its behavior still oscillates around the ground truth, and its validation loss, while better than the one of LTCs, remains relatively large, as shown in Figure 2.
3 Liquid Resistance Liquid Capacitance Networks (LRCs)
Definition. For simplicity, LTCs and STCs assume that the membrane capacitance is constant and equal to one. This is however not the case in biological neurons. They have a nonlinear dependence on the MPs and the input (Howell et al., 2015; Severin et al., 2022; Kumar et al., 2023). LRCs take this aspect into account, by adding a liquid elastance (reciprocal of capacitance) $\epsilon(w_{i})$, in Equation (2), where the weights matrix $o$ and the bias vector $p$ play the same role as the ones in an artificial neuron:
$\begin{array}[]{c}\dot{h}_{i}=\epsilon(w_{i})\,(-\sigma(f_{i})\,h_{i}+\tau(u_{i})\,e_{li})\\[5.69054pt] w_{i}=\sum_{j=1}^{m+n}o_{ji}y_{j}+p_{i}\end{array}$ (3)
Thus, the time constant in LRCs factors into a liquid resistance and a liquid capacitance. The second, acts as an additional control of the time constant, by increasing it when the desired behavior is smooth, and leaving it unchanged when the desired behavior is stiff. For convenience, we provide the definition of both an asymmetric and a symmetric form of elastance in Equations (4-5), respectively. Here, the parameters vector $k$ of the symmetric version is non-negative in each of its components:
$(A)\quad$ $\epsilon(w_{i})=\sigma(w_{i})$ (4)
$(S)\quad$ $\epsilon(w_{i})=\sigma(w_{i}+k_{i})-\sigma(w_{i}-k_{i})$ (5)
Figure 4: Left, Middle: The asymmetric and symmetric shape of the membrane’s elastance $\epsilon(w_{i})$ proposed for LRCs, using Formulas (4) and (5), respectively. Right: Experiment showing that the LCR’s elastance dampens the dynamic behavior of STCs in an input and state-dependent way. It can fully reduce it to zero, or keep it as it is, indicated by the red and black dashed lines, respectively.
@@ 3 Liquid Resistance Liquid Capacitance Networks (LRCs)
Smoothness. The dynamic behavior of the LRC output neuron in the Lane-Keeping Task is shown in Figure 1(right). It has a very smooth behavior, closely following the ground truth.
Due to its very smooth behavior, an explicit Euler integration with one unfolding, is often sufficient for obtaining an acceptable validation loss. This dramatically speeds up computation, as the solver introduces no overhead. This is particularly useful for time series with regularly sampled data, lacking timing information. Here, a time step of one, leads to a network closely resembling gated recurrent neural networks (Hochreiter and Schmidhuber, 1997; Cho et al., 2014; Zhou et al., 2016). Its architecture is shown in Figure 3 (right). We call this network, an LRC unit (LRCU). For further details and discussion on this, see Appendix A.2. As we show in the experimental-results section in Section 4.1, LRCUs outperform gated RNNs on the RNN benchmarks, when used wih either a symmetrical elastance (LRCU-S) or an asymmetrical elastance (LRCU-A).
Theorem 1 (LRC Smoothness).
@@
The proof of Theorem 1 is given in the Appendix A.1.1. Intuitively, for an LRC having the parameter values for its forget and update conductances fixed as in an STC, one can set the parameter values of its elastance, such that the Lipschitz constant of the LRC is smaller than the one of the STC. This reduces the oscillations not only on the training data, but also on the validation (and testing) data.
@@
Accuracy. The addition of a liquid elastance in LRCs does not only reduce oscillations, but it also dramatically improves accuracy. This is illustrated for the Lane-Keeping Task in Figure 2, where the validation loss of the LRC’s output neuron, is considerably smaller compared to the one of LTCs and STCs. While the above theorem proves smoothness, the following theorem proves accuracy.
Theorem 2 (LRC Accuracy).
@@
Let $h^{s}$ be an STC-model instance of Equation (2) with $h^{r}$ being an associated LRC-model instance of Equation (3) with $\lambda^{r}<\lambda^{s}$ (Theorem 1). Let $o^{T}_{t}$ be training labels and $\hat{o}^{T,r}_{t}$ and $\hat{o}^{T,s}_{t}$ be the predictions of the models $h^{r}$ and $h^{s}$ respectively. Let both models have a small training loss $\|o^{T}_{t}-\hat{o}^{T}_{t}\|_{2}<\epsilon_{T}$ with small $\epsilon_{T}>0$. Let $y^{T}_{t}$ and $y^{V}_{t}$ be respectively training and validation input. Let the validation set be similar to the training set in a sense that for every time $t$ it holds that $\|y^{T}_{t}-y^{V}_{t}\|_{2}<\epsilon_{y}$ and $\|o^{T}_{t}-o^{V}_{t}\|_{2}<\epsilon_{o}$ with small $\epsilon_{y},\epsilon_{o}>0$. Then it holds that the upper bound of the validation loss of the LRC-model is smaller than the one of the STC-model.
@@
The proof of Theorem 2 is given in the Appendix A.1.2. Intuitively, if the validation set is similar to the training set modulo some small perturbation, one can use a Taylor series approximation of the validation loss $o(y^{V})\approx o(y^{T})+(y^{V}-y^{T})^{T}\nabla_{y}o(y^{V})$ and get an upper bound on the validation loss by using Theorem 1. As the Lipschitz constant of the LRC is smaller than the one of STC, also the upper bound of the validation loss of the LRC-model is smaller than the one of the STC-model.
@@
Efficiency. Solving LTCs, STCs, and LRCs with popular ODE-integration techniques, is computationally expensive. A hybrid forward-backward Euler method (Lechner et al., 2020) for example, takes multiple unfolding rounds for a given input, to achieve a good approximation of the solution. In Figure 2 we show the validation loss of LTCs, STCs, and LRCs, for the Lane-Keeping Task (Lechner et al., 2022), where the state of the neurons is computed with a fixed integration step $\Delta$. Diving this into 6 equidistant time steps $\Delta{/}6$ within one update step, results in a considerably better model, but at a higher cost of 10min versus 4min per epoch. This results in a 2.5×2.5$\times$2.5 × longer computation time.
@@
Interpretability. In Lechner et al. (2020), it was shown that trained LTCs are interpretable, when used in a neural-circuit-policy (NCP) architecture. Later on in Farsang et al. (2023), it was shown that interpretability naturally occurs in STCs, even when using a single, dense, all-to-all hidden layer. This is all the more true for LRCs in general, and LRCUs in particular. As we show in Section 4 for the Lane-Keeping Task, LRCUs and more so LRCs solved with more sophisticated solvers, lead to very interpretable behavior of the command neurons (those impacting the output) of the hidden layer.
4 Experimental Results
In the previous sections we defined LRCs as an extension of STCs, by arguing that using a constant capacitance, as is the case in LTCs and STCs, is not only biologically implausible, but it also leads to large oscillations in the trained networks. We also showed that by considering a liquid (that is a state-and-input dependent) capacitance, dramatically reduces the validation loss of the resulting network, by smoothing out the oscillations. Due to smoothness, solving LRCs with explicit, one-unfolding Euler, leads to very small validation loss, without any computational overhead. We called LRCs integrated this way, LRCUs, as they closely resemble gated RNNs when using a time-step of one.
In the following subsections, we present our experimental results for LRCs, on three different topics of interest in the ML community. The first investigates the accuracy and speed of LRCUs in comparison with the one of LSTMs, GRUs, and MGUs on classic benchmarks used for RNNs. The second investigates the accuracy and speed of LRCs in solving classic Neural-ODE tasks. Finally, the third investigates the interpretability of the learned LRCUs (and thus LRCs) for the Lane-Keeping Task.
4.1 Accuracy and Speed of LRCUs on RNNs Benchmarks

### v4

@@ Abstract
We introduce liquid-resistance liquid-capacitance neural networks (LRCs), a neural-ODE model which considerably improves the generalization, accuracy, and biological plausibility of electrical equivalent circuits (EECs), liquid time-constant networks (LTCs), and saturated liquid time-constant networks (STCs), respectively. We also introduce LRC units (LRCUs), as a very efficient and accurate gated RNN-model, which results from solving LRCs with an explicit Euler scheme using just one unfolding. We empirically show and formally prove that the liquid capacitance of LRCs considerably dampens the oscillations of LTCs and STCs, while at the same time dramatically increasing accuracy even for cheap solvers. We experimentally demonstrate that LRCs are a highly competitive alternative to popular neural ODEs and gated RNNs in terms of accuracy, efficiency, and interpretability, on classic time-series benchmarks and a complex autonomous-driving lane-keeping task.
Keywords:
Machine Learning Bio-inspired Neural Networks Liquid Capacitance.
@@ 1 Introduction
Electrical equivalent circuits (EECs) are the foremost mathematical model used in neuroscience for capturing the dynamic behavior of biological neurons [17, 36]. EECs associate the membrane of a neuron with a capacitor, whose potential varies according to the sum of the stimulus, leaking, and synaptic currents, passing through the membrane. Consequently, EECs are formulated as a set of capacitor ordinary differential equations (ODEs).
@@
EECs, however, received little attention in the ML community due to their ODE nature, except for continuous-time recurrent neural networks (CT-RNNs) [13], which are arguably EECs with electrical synapses [12]. This changed with the introduction of Neural ODEs [6] and later their latent [29] and augmented [11] versions. EECs with chemical synapses are called liquid time-constant networks (LTCs) [14], as their time constant depends on both the state and the input.
@@
The main promise of LTCs and Neural ODEs is to better capture physical reality and bridge the gap between natural sciences and ML. They are able to recover missing data in irregularly-sampled time series, which was investigated in LTCs [14], the GRU-ODE-Bayes method [5], Neural Controlled Differential Equations [18] and the Neural Lad framework [26]. Furthermore, LTCs were shown to be more interpretable in an autonomous lane-keeping task both in real-world [25] and simulation environments [12]. An important obstacle in the use of Neural ODEs and especially LTCs, however, is their stiff oscillatory nature, which requires the use of expensive ODE solvers [3].
@@
In this paper, we show that the stiff dynamic behavior of LTCs results from abstraction. For simplicity, EECs and LTCs first ignore the saturation aspects of membrane’s ionic channels [1], as discussed in [12]. To take such aspects into account, they introduce saturated LTCs (called STCs). Second, and more importantly, EECs, LTCs and STCs also assume that membranal capacitance is constant, which is disproved by recent results demonstrating a nonlinear dependence on the neural state and input [16, 30, 20].
@@
Here, we show that adding a liquid membrane capacitance to STCs, considerably reduces their variation with respect to inputs, while also enhancing their accuracy. We call the resulting model a liquid-resistance liquid-capacitance neural network (LRC), as the LRC’s liquid time-constant is now factored this way. Due to the gentle varying LRC behavior, an explicit Euler integration scheme with one unfolding, is often enough to obtain a very small error. We call LRCs solved with this scheme, LRC units (LRCUs). They closely resemble and improve on the performance of gated recurrent units, such as Long Short-Term Memory (LSTMs) [15], Gated Recurrent Units (GRUs) [8] and Minimal Gated Units (MGUs) [37]. For irregularly sampled time series, however, one can use more expensive solvers with LRCs, too.
@@
We experimentally evaluate LRCs with respect to their accuracy and speed on the classic gated-RNN and Neural-ODE benchmarks, respectively, and with respect to their interpretability on a complex autonomous-driving lane-keeping task. Our results show that LRCs outperform LSTMs, GRUs, and MGUs on RNN benchmarks, are able to successfully solve Neural-ODE tasks, and at the same time, are much more interpretable than LSTMs, GRUs, and MGUs on the lane-keeping task.
@@
  • We introduce Liquid-Resistance Liquid-Capacitance Networks (LRCs), which considerably improve the generalization, accuracy, and biological plausibility of EEC, LTC, and STC models, respectively.
@@
  • We prove that the liquid capacitance of LRCs, leads to a more gentle varying behavior than in LTCs and STCs, while at the same time dramatically increasing their accuracy even for cheap solvers.
@@
The rest of this paper is organized as follows. In Section 2, we review EECs, LTCs, and STCs. Next, we introduce Liquid-Resistance Liquid-Capacitance Networks (LRCs) and LRCUs and discuss and prove their properties. In Section 4 we provide our very encouraging experimental results. Finally, in Section 5, we draw our conclusions and discuss future work.
@@ 2.1 Liquid Time-Constant Networks (LTCs)
Definition. EECs are a simplified electrical model, defining the dynamic behavior of the membrane-potential (MP) of a post-synaptic neuron, as a function of the MP of its pre-synaptic neurons, for electrical or chemical synapses [17, 36]. EECs with chemical synapses only, are also known in the ML community as LTCs [24, 14, 25]. An LTC with $m$ neurons and $n$ inputs is defined as follows:
[…]
It states that the rate of change of the MP $h_{i}$ of neuron $i$, is the sum of its forget current $f_{i}\,h_{i}$ and its update current $u_{i}\,e_{li}$. The forget conductance $f_{i}$ is the liquid time-constant reciprocal of $h_{i}$, which depends on y=[h,x]y\,{=}\,[h,x], where $h$ represents the MPs of presynaptic neurons and $x$ is the external inputs. Here, $g_{ji}$ is the maximum conductance of the synaptic channels, $a_{ji},b_{ji}$ are parameters controlling the sigmoidal probability of these channels to be open, and $g_{li}$, $e_{li}$ are the membrane’s leaking (resting) conductance and potential, respectively. In the update conductance $u_{i}$, parameters $k_{ji}\,{=}\,g_{ji}e_{ji}{/}e_{li}$ take the sign of $e_{ji}{/}e_{li}$, where $e_{ji}$ is the synaptic channel’s reversal potential, that is, the MP at which there is no net ionic flow.
@@ 2.2 Saturated Liquid Time-Constant Networks (STCs)
Definition. For simplicity, LTCs ignore saturation in synaptic channels [1]. One can take this aspect into account, by constraining the forget and signed-update conductances $f_{i}$ and $u_{i}$ with a sigmoid and a hyperbolic tangent, to range within [0,1] and [-1,1], respectively. Letting $f_{i}$ and $u_{i}$ be as above, this leads to the definition of saturated LTCs (STCs) of [12]:
$\begin{array}[]{c}\dot{h}_{i}=-\sigma(f_{i})\,h_{i}+\tau(u_{i})\,e_{li}.\end{array}$ (2)
3 Liquid-Resistance Liquid-Capacitance Networks (LRCs)
3.1 Model Definition
Definition. LTCs and STCs assume that the membrane capacitance is constant and equal to one. This is however not the case in biological neurons. They have a nonlinear dependence on the MPs and the incoming inputs [16, 30, 20]. LRCs take this aspect into account, by adding a liquid elastance (reciprocal of capacitance) $\epsilon(w_{i})$, in Equation (2), where the weights matrix $o$ and the bias vector $p$ play the same role as the ones in an artificial neuron:
$\begin{array}[]{c}\dot{h}_{i}=\epsilon(w_{i})\,(-\sigma(f_{i})\,h_{i}+\tau(u_{i})\,e_{li})\\[5.69054pt] w_{i}=\sum_{j=1}^{m+n}o_{ji}y_{j}+p_{i}.\end{array}$ (3)
Thus, the time constant in LRCs factors into a liquid-resistance and a liquid-capacitance. The second, acts as an additional control of the time constant, by increasing it when the desired variation of the function is gentle, and leaving it unchanged when the desired behavior is stiff. For convenience, we provide the definition of both an asymmetric and a symmetric form of elastance in Equations (4-5), respectively. Here, the parameters vector $k$ of the symmetric version is non-negative in each of its components:
$(A)\quad$ $\epsilon(w_{i})=\sigma(w_{i}),$ (4)
$(S)\quad$ $\epsilon(w_{i})=\sigma(w_{i}+k_{i})-\sigma(w_{i}-k_{i}).$ (5)
Refer to captionRefer to caption
Figure 1: The asymmetric and symmetric shape of the membrane’s elastance $\epsilon(w_{i})$ proposed for LRCs, using Formulas (4) and (5), respectively.
@@ 3.1 Model Definition
3.2 Model Properties
In this section, we investigate important model properties of LRCs. For this purpose, we focus on the complex autonomous driving Lane-Keeping experiment, as it is a regression task where LTCs have been shown to perform well. We aim to determine if we can improve upon the properties of LTCs by incorporating the additional saturation aspect and liquid capacitance of LRCs. Specifically, we are interested in examining the Lipschitz constant, model accuracy, and efficiency.
Figure 3: Dynamic behavior of the output of LTC, STC and LRC in the regression task of Lane-Keeping [23].
Figure 4: Experiment showing that the LRC’s elastance dampens the dynamic behavior of STCs in an input and state-dependent way. It can fully reduce it to zero, or keep it as it is, indicated by the red and black dashed lines, respectively.
Lipschitz constant.
In the case of LTCs, shown on the left in Figure 3, the output has an oscillatory behavior, with a relatively large validation loss. Moreover, LTCs fail to converge with an explicit Euler integration scheme of order one and just one unfolding. In contrast, STCs convergence with an explicit Euler integration scheme of order one and just one unfolding presented in the middle figure. However, the output neuron still exhibits stiff oscillatory behavior. In LRCs, this neuron has a very gentle varying behavior, with a very small validation loss, even when using an explicit Euler scheme of order one and just one unfolding (no overhead). This highly advantageous, as it was demonstrated that a lower Lipschitz constant enhances robustness [9], generalization [32], and interpretability [33].
Due to its very gentle varying behavior of LRCs, an explicit Euler integration with one unfolding, is often sufficient for obtaining an acceptable validation loss. This dramatically speeds up computation, as the solver introduces no overhead. This is particularly useful for time series with regularly sampled data, lacking timing information. Here, a time step of one, leads to a network closely resembling gated recurrent neural networks [15, 8, 37]. Its architecture is shown in Figure 2 (right). We call this network, an LRC unit (LRCU). For further details and discussion on this, see Appendix 0.A.2. As we show in the experimental-results section in Section 4.1, LRCUs outperform gated RNNs on the RNN benchmarks, when used wih either a symmetrical elastance (LRCU-S) or an asymmetrical elastance (LRCU-A).
Theorem 3.1 (LRC Lipschitz).
@@ 3.2 Model Properties
The proof of Theorem 3.1 is given in the Appendix. Intuitively, for an LRC having the parameter values for its forget and update conductances fixed as in an STC, one can set the parameter values of its elastance, such that the Lipschitz constant of the LRC is smaller than the one of the STC. This leads to LRC being a function which does not change rapidly when the input changes.
Accuracy and Generalization.
@@
The addition of a liquid elastance in LRCs does not only reduce oscillations, but it also dramatically improves accuracy. This is illustrated for the Lane-Keeping Task in Figure 5, where the validation loss of the LRC’s output neuron, is considerably smaller compared to the one of LTCs and STCs. While the above theorem proves a smaller Lipschitz constant, the following theorem proves generalization.
Figure 5: Training loss in the first and validation loss in the second figure of LTCs, STCs, and LRCs for the Lane-Keeping Task.
Theorem 3.2 (LRC Generalization).
@@
Let $h^{s}$ be an STC-model instance of Equation (2) with $h^{r}$ being an associated LRC-model instance of Equation (3) with $\lambda^{r}<\lambda^{s}$ (Theorem 3.1). Let $o^{T}_{t}$ be training labels and $\hat{o}^{T,r}_{t}$ and $\hat{o}^{T,s}_{t}$ be the predictions of the models $h^{r}$ and $h^{s}$ respectively. Let both models have a small training loss $\|o^{T}_{t}-\hat{o}^{T}_{t}\|_{2}<\epsilon_{T}$ with small $\epsilon_{T}>0$. Let $y^{T}_{t}$ and $y^{V}_{t}$ be respectively training and validation input. Let the validation set be similar to the training set (drawn in-distribution) in a sense that for every time $t$ it holds that $\|y^{T}_{t}-y^{V}_{t}\|_{2}<\epsilon_{y}$ and $\|o^{T}_{t}-o^{V}_{t}\|_{2}<\epsilon_{o}$ with small $\epsilon_{y},\epsilon_{o}>0$. Then it holds that the upper bound of the validation loss of the LRC-model is smaller than the one of the STC-model.
@@
The proof of Theorem 3.2 is given in the Appendix. Intuitively, if the validation set is similar to the training set modulo some small perturbation, are drawn from the same distribution, one can use a Taylor series approximation of the validation loss $o(y^{V})\approx o(y^{T})+(y^{V}-y^{T})^{T}\nabla_{y}o(y^{V})$ and get an upper bound on the validation loss by using Theorem 3.1. As the Lipschitz constant of the LRC is smaller than the one of STC, also the upper bound of the validation loss of the LRC-model is smaller than the one of the STC-model.
Efficiency.
@@
Solving LTCs, STCs, and LRCs with popular ODE-integration techniques, is computationally expensive. A hybrid Euler method [25] for example, takes multiple unfolding rounds for a given input, to achieve a good approximation of the solution. In Figure 5 we show the validation loss of LTCs, STCs, and LRCs, for the Lane-Keeping Task [23], where the state of the neurons is computed with a fixed integration step $\Delta$. Diving this into 6 equidistant time steps $\Delta{/}6$ within one update step, results in a considerably better model, but at a higher cost of 10min versus 4min per epoch. This results in a $2.5\times$ longer computation time.
Table 1: Training details of LTC, STCs and LRCs on the Lane Keeping dataset.
Model Time Per Epoch Best Epoch Lipschitz Constant
LTC H6 10.06 min $52\pm 06$ $90{\pm}23{\cdot}10^{-4}$
STC E1 03.97 min $50\pm 22$ $101{\pm}25{\cdot}10^{-4}$
STC H6 10.13 min $69\pm 10$ $89{\pm}23{\cdot}10^{-4}$
LRC-A E1 04.05 min $43\pm 08$ $47{\pm}08{\cdot}10^{-4}$
LRC-S E1 04.16 min $22\pm 03$ $43{\pm}11{\cdot}10^{-4}$
Table 1 presents that per epoch and curves, the training took around 10 minutes for 6 unfoldings per step and 4 minutes for 1 unfolding per step, which is 2.5 times faster. As the curves of LTCs and STCs with the hybrid solver converge in around 50-70 epochs, and the LRC-S and LRC-A in around 20 and 40, respectively, the total speedup until the best epoch of the latter with respect to the former is 3-7.8 times, while also being a much more accurate. Note that solving STCs with Explicit Euler and one unfolding is computationally inexpensive, but its validation loss is not very satisfactory.
@@
4 Further Experimental Results
In the previous sections we defined LRCs as an extension of STCs, by arguing that using a constant capacitance, as is the case in LTCs and STCs, is not only biologically implausible, but it also leads to large oscillations in the trained networks. We also showed that by considering a liquid (that is a state-and-input dependent) capacitance, dramatically reduces the validation loss of the resulting network, by having better generalization. Due to the more gently varying behavior, solving LRCs with explicit, one-unfolding Euler, leads to very small validation loss, without any computational overhead. We called LRCs integrated this way, LRCUs, as they closely resemble gated RNNs when using a time-step of one.
In the following subsections, we present our further experimental results for LRCs, on three different topics of interest in the ML community. The first investigates LRCUs in comparison with the one of LSTMs, GRUs, and MGUs on classic benchmarks used for RNNs. The second investigates LRCs in solving classic Neural-ODE tasks. Finally, the third investigates the interpretability of the learned LRCUs (and thus LRCs) for the Lane-Keeping Task.
4.1 LRCUs on RNNs Benchmarks